# ViGoR: Improving Visual Grounding of Large Vision Language Models with Fine-Grained Reward Modeling

### v1

@@ 4 Experiments and Results
In this section, we analyze the effectiveness of our approach in comparison with competitive baselines, and delve into the contributions from each component in our design. Finally, we provide visualizations to qualitatively compare the output of fine-tuned model with that of its initial state. In the following, “ViGoR-AT” refers to our approach utilizing only reward modeling with automatic methods. “ViGoR-RM” represents our method employing reward modeling with fine-grained human feedback. “ViGoR-All” and “ViGoR” denotes our method with the combination of the reward modeling with automatic methods and fine-grained human feedback.
@@ 4.1 Testing Framework
With computation efficiency in mind, we select the smallest variant of the Grounding DINO model with the Swin-T [22] backbone with the official checkpoint and the default box and text thresholds for the open-set object detection task. Our proposed method can likely directly benefit from the larger (but more computationally expensive) variants or future advances in these models.
@@ 4.2 LVLM Model Fine-tuning
To support our experiment, we used 25,574 images from the ADE20K dataset [34] (training split) as the basis, as they cover a wide variety of relatively complex scenes. To preserve the generalization of our method, we disregard the original annotations of the dataset. We elicit responses from our base LVLM through a set of question prompts (e.g. “Please provide a detailed description of the given image.”), and sample five different outputs for each input image. Subsequently, these outputs are fed into our reward model to obtain a corresponding reward score for each description. In addition to keeping the highest-scoring description following the common practice in rejection sampling, we enhance this description further by applying a refinement module to achieve additional quality improvements.
@@
ViGoR-All 1.98 1.96 2.00 1.98 1.99 1.97 1.97 1.98
Table 1: Detailed caption analysis via GPT-4Vision ranking. Scores are the average preference rank of the caption when compared within the 4 candidates in the table (ranging from 1 to 4; lower is better).
@@
4.3 Reward Model Training
Human feedback collection.
To allow the reward model to be highly receptive to errors in the descriptions, we use a pre-trained checkpoint for LLaVA that is already fine-tuned with our automatic reward generation mechanism. Through hands-on experimentation, we observed that while the images from the ADE20K dataset exhibit excellent scene variability, the resulting cognitive load for human annotators is very high. As such, instead, we generate 16,000 detailed image captions using images selected from the somewhat simpler MS COCO [19] dataset. We enlist the services of a team of 15 professional annotation workers to assist us in the creation of the evaluation data set for the training of the reward model using the process described in Section 3.1. We provide carefully designed and detailed annotation instructions (available in our Supplementary Materials) to our annotators, along with extensive sample annotations. The process took approximately 3 weeks.
Reward model.
We initialize the same model architecture as our base LVLM with the same weights used to generate the annotation samples as the starting state of our reward model. We train this model for 5 epochs on the dataset with 16,000 samples with a batch size of 32 and an initial learning rate of $2\times 10^{-5}$.
Method #P VQA Random Popular Adversarial
@@ 4.3 Reward Model Training
mPLUG-Owl 7B 53.3 51.7 50.6 50.3 50.7 50.3
MiniGPT-4 7B 77.8 75.4 68.3 64.3 66.6 62.45
MM-GPT 9B 50.0 50.0 50.0 50.0 50.0 50.0
InstructBLIP 13B 88.7 85.0 81.4 75.1 74.4 67.7
LLaVA 7B 54.4 52.3 52.4 51.3 50.8 50.4
ViGoR (ours) 7B 85.1 89.0 81.5 83.0 75.5 73.8
Improvement +30.7 +36.7 +29.1 +31.7 +24.7 +23.4
Table 2: Quantitative results on POPE benchmark.
4.4 Quantitative Results
The types of natural interactions that a user can have with a LVLM are essentially unlimited in variety. However, since the focus of our work is on improving visual grounding, we primarily evaluate the benefits of our technique in two types of settings. First, we analyze the unstructured output of the LVLM when it is asked to generate a detailed and comprehensive caption for an image. This tests the model’s ability to be perceptive to all regions and objects in the image. Second, we evaluate the model’s performance on responding to targeted questions for which a concise and unambiguous answer is required. This further tests the model’s ability to allow the input prompt to guide its visual understanding and formulate an accurate response.
Method #Param Data VQA Data EX CT PO CO PT CE SC LM AT OC Overall
@@ 4.4 Quantitative Results
LLaVA [^1] We report the number of LLaVA v1.0 reproduced by ourselves. 7B 158K 158.3 83.3 51.7 85.0 94.2 85.0 145.8 125.8 74.3 57.5 960.1
@@
Table 3: Quantitative results on MME benchmark. “VQA Data” indicates whether the model is fine-tuned using the VQA datasets. For a fair comparison, we compare the Large Vision Language Model using the Vicuna backbone.
Method HL CA CO MA Overall
InstructBLIP 187.98 110.20 143.42 104.14 545.74
LLaVA 156.19 61.44 97.84 62.26 377.73
ViGoR-AT 183.81 106.34 124.36 122.08 536.59
ViGoR-RM 178.92 94.89 133.33 137.82 544.96
ViGoR-All 187.61 112.08 135.22 140.40 575.31
Table 4: Quantitative results on our MMViG benchmark.
4.4.1 GPT4-Assistant Evaluation
Objective quantitative measurement of free-form responses has been a major challenge in the LLM and LVLM space due to the inherent ambiguities in defining quality. Furthermore, it is unclear how structured information can be extracted from generated text so that a score can be assigned using traditional objective functions. To address this, we utilize the advanced capabilities of the GPT-4Vision service to evaluate the outputs. We ask GPT-4Vision to rank the output of different LVLM candidates according to numerous criteria that signify the quality of the captions. This offers a detailed view of the visual foundation and text generation capabilities of the LVLMs under test. We perform this evaluation with 96 randomly selected images from the COCO validation dataset and show the results in Table 1 along with an overview of the definitions of the metrics. We select the four most relevant configurations for comparison: 1) the original LLaVA baseline model that we use as the initialization; 2) ViGoR-AT: the ViGoR method using only reward modeling with automatic methods; 3) ViGoR-RM: the ViGoR method using only the guidance from the reward model trained using the annotator feedback, and lastly 3) ViGoR-All: the complete model. As is clearly evident, our complete model consistently achieves the best ranking within the comparison across all metrics. Furthermore, we observe that either source of supervision signals significantly outperforms the original baseline model. Further details are provided in the supplementary materials.
@@ 4.4.1 GPT4-Assistant Evaluation
4.4.2 Short Answer Benchmark Evaluation
Numerous benchmarks have been released recently to support the development of LVLMs. To decrease the complexity and ambiguity, these benchmarks typically consist of questions with concise answers such as “Is there a red apple in the image?” As this style of questioning is not explicitly considered in our model training, these benchmarks measure both the visual grounding capabilities and the generalization of our method. As such, we select the commonly used POPE [18] and MME [11] benchmarks, which are constructed using images from the validation set of the COCO dataset [19]. POPE probes the LVLM with 3000 yes/no questions targeting 500 images to evaluate the LVLM’s ability to determine the existence of specific objects in a scene. MME has a similar questioning setup, but extends the questioning subject to cover aspects such as numerical count, position, and color of objects, as well as other vision tasks such as scene/landmark identification and OCR.
As illustrated in Tables 2 and  3, our model shows consistent improvement across all categories evaluated in both datasets. Furthermore, while the competitive baseline methods use VQA datasets containing ground truth for these types of straightforward questions, our method is able to achieve comparable or superior results without fine-tuning on such resources. This suggests that the visual grounding abilities learned by the model through our process is general. When the model learns to generate accurate comprehensive descriptions for images, it is able to apply the capabilities to related reasoning tasks, suggesting that the logical deduction capabilities of the underlying pre-trained LLM is retained and leveraged. Furthermore, our 7 billion parameter model shows on par performance with some models with nearly double the number of parameters (13 billion). This underscores the efficiency and robustness of our model in handling diverse visual-language tasks.
However, POPE only measures object-level hallucinations or omissions, while MME provides as few as 60 question/answer ground truth pairs over 30 images for the various categories. To increase the comprehensiveness of our analysis, we introduce a new benchmark, MMViG, to assess the visual grounding capabilities of LVLMs in four distinct categories: Hallucinations (HL), Counting Accuracy (CA), Color Accuracy (CO), Material Accuracy (MA). Our evaluation process involves selecting 200 images from the MS COCO 2017 validation data set for each category. We create a set of 400 paired yes/no questions across the four categories. We automatically generate the questions for HL, CA, CO and MA using ground-truth labels from both the COCO and PACO [26] annotations. This methodology allows us to thoroughly examine the LVLM’s ability to accurately interpret and respond to a diverse range of visual inputs. As shown in Table 4, the improvements and trends observed in previous scenarios are also evident in these assessments.
@@ 4.5 Qualitative Results
In Figure 3, we show qualitative results of our technique in action, focusing on a comparative analysis between our model and its starting point (LLaVA). This comparison is particularly relevant as our model undergoes fine-tuning based on the LLaVA model. By juxtaposing these two models, we highlight the enhancements and advancements our model has achieved, offering valuable insights into the efficacy and impact of our training strategies.
AT-RM HF-RM Refine MMViG
- - 521.41
- 536.59
- 544.96
- 551.87
575.31
Table 5: Ablation studies on different components of our framework.
4.6 Ablation Studies
In this section, we present ablation studies of our approach to justify various design choices. For simplicity, we evaluated all variants on the MMViG benchmark.
4.6.1 Ablation Studies of Framework Components
reward model design MMViG
none 377.73
holistic-based 403.01
fine-grained-based 544.96
@@ 4.6.1 Ablation Studies of Framework Components
In this section, we examine the efficacy of different components within our framework. Table 5 presents the performance of each module, underscoring their individual significance. The results clearly demonstrate that integrating all components yields the most optimal performance.
4.6.2 Fine-grained vs Holistic Evaluation by Reward Model
In this section, we perform a comparative assessment between the fine-grained analysis of sentences by the reward model as compared with the more commonly used holistic evaluation of complete responses. This assessment is carried out in the context of reward modeling with human feedback. For the holistic reward model, we retrained a reward model using a singular, unified reward for each description. As shown in Table 6, the fine-grain-based approach demonstrates better performance than the holistic approach, which shows that denser and more informative signals allow the reward model to better discern the quality of the generated text by creating more direct links between the visual features in the image and the sentences describing them.
5 Conclusion and Future Work
@@
In the future, we plan to apply RLHF training with our ViGoR framework, which may offer additional improvements over our training based on rejection sampling. As well, we anticipate that explicit visual alignment predictions through reasoning about object and element localization can further improve the visual grounding of LVLMs.

### v3

@@ 4 Experiments and Results
We compare the effectiveness of our approach with competitive baselines, and delve into the contributions from each component in our design. As well, we provide visualizations to qualitatively compare the output of fine-tuned model with that of its initial state. In the following, “ViGoR-AT" refers to our approach utilizing only reward modeling with automatic methods. “ViGoR-RM" represents our method employing reward modeling with fine-grained human feedback. “ViGoR-All" and “ViGoR" denote our method with the combination of the reward modeling with automatic methods and fine-grained human feedback.
@@ 4.1 Testing Framework
With computation efficiency in mind, we select the smallest variant of the Grounding DINO model with the Swin-T [24] backbone with the official checkpoint and the default box and text thresholds (0.25) for the open-set object detection task. Our proposed method can likely directly benefit from the larger and more computationally expensive variants or future advances in these models.
@@ 4.2 LVLM Model Fine-tuning
To support our experiment, we used 25,574 images from the ADE20K dataset [37] (training split) as the basis, as they cover a wide variety of relatively complex scenes. To preserve the generalization of our method, we disregard the original annotations of the dataset. We elicit responses from our base LVLM through a set of question prompts (e.g. “Please provide a detailed description of the given image."), and sample 5 outputs for each input image. Subsequently, these outputs are fed into our reward model to obtain a corresponding reward score for each description. In addition to keeping the highest-scoring description following the common practice in rejection sampling, we enhance this description further by applying a refinement module to achieve additional quality improvements.
@@
4.3 Reward Model Training
Human feedback collection.
To allow the reward model to be highly receptive to errors in the descriptions, we use a pre-trained checkpoint for LLaVA that is already fine-tuned with our automatic reward generation mechanism. Through hands-on experimentation, we observed that while the images from the ADE20K dataset exhibit excellent scene variability, the resulting cognitive load for human annotators is very high. As such, instead, we generate 15,440 detailed image captions using images selected from the somewhat simpler MS COCO [21] dataset. We enlist the services of 15 professional annotators to assist us in creating the evaluation data set for the training of the reward model using the process described in Section 3.1. We provide carefully designed and detailed annotation instructions (available in our Supplementary Materials) to our annotators, along with extensive sample annotations. The process took approximately 3 weeks.
Reward model.
We initialize the same model architecture as our base LVLM with the same weights used to generate the annotation samples as the starting state of our reward model. We train this model for 5 epochs on the dataset with 15,440 samples with a batch size of 32 and an initial learning rate of $2\times 10^{-5}$.
@@ 4.3 Reward Model Training
ViGoR-All 1.97 1.96 1.99 1.98 1.99 1.97 1.97 1.97
Table 1: Detailed description generation evaluation. Scores are the average preference rank of the caption when compared within the 4 candidates in the table as judged by GPT-4Vision (ranging from 1 to 4; lower is better).
@@
4.4 Quantitative Results
The types of natural interactions that a user can have with a LVLM are essentially unlimited in variety. Since the focus of our work is on improving visual grounding, we primarily evaluate the benefits of our technique in three types of settings. First, we analyze the unstructured output of the LVLM when it is asked to generate a detailed and comprehensive caption for an image. This tests the model’s ability to be perceptive to all regions and objects in the image. Second, we evaluate the ability of the model to form short sentences in response to targeted questions where some logical reasoning is needed. Lastly, we evaluate the model’s performance on responding to highly specific questions for which a concise and unambiguous answer is required. This further tests the model’s ability to use the input prompt to guide its visual understanding. We submit that this variety of tasks forms a comprehensive testing framework.
4.4.1 Detailed Description Generation Evaluation.
Examining the factuality of detailed descriptions generated by a LVLM based on an input image allows provides highly comprehensive insight into visual grounding capabilities. However, objective quantitative measurement of free-form responses has been a major challenge in the LLM and LVLM space due to the inherent ambiguities in defining quality. Furthermore, it is unclear how structured information can be extracted from generated text so that a score can be assigned using traditional objective functions. As demonstrated by [36], the GPT family of models have very high levels of agreement with human annotators when asked to follow instructions to evaluate output of other models. This is likely due to the large quantities of proprietary data used for training. Therefore, we ask GPT-4Vision to rank the output of different LVLM candidates according to numerous criteria that signify the quality of the captions. This offers a detailed view of the visual foundation and text generation capabilities of the LVLMs under test.
For detailed insight, we evaluate model output in terms of hallucinations, relevance, reasoning, level of detail, as well as accuracy in counting, attributes, and relationships between objects. As it is difficult to grade lengthy text responses against an absolute scale, we instead use preference ranking. GPT-4Vision is asked to consider several candidate responses relative to one another against the sample input image and text query, and provide a ranked list over the responses for each metric. The rankings are averaged over all samples to obtain an overall ranking score for each model, where lower is better.
We evaluate using 300 randomly selected images from the COCO validation dataset and show the results in Table 1 along with an overview of the definitions of the metrics. We select the four most relevant configurations for comparison: 1) the original LLaVA baseline model that we use as the initialization; 2) ViGoR-AT: the ViGoR method using only reward modeling with automatic methods; 3) ViGoR-RM: the ViGoR method using only the guidance from the reward model trained using the annotator feedback, and lastly 3) ViGoR-All: the complete model. As is clearly evident, our complete model consistently achieves the best ranking within the comparison across all metrics. Furthermore, we observe that either source of supervision signals significantly outperforms the original baseline model. Further details are provided in the Supplementary Materials.
Method #Param MMHal-Bench
LLaVA 7B 1.3
LLaVA-RLHF 7B 1.4 (+0.1)
ViGoR (ours) 7B 1.6 (+0.3)
Table 2: Short answer evaluation on MMHal-Bench benchmark. For a fair comparison, we show the results of LLaVA-RLHF using solely RLHF optimization without supervised fine-tuning using additional data.
4.4.2 Short Answer MMHal-Bench Evaluation.
Next, we showcase the performance of our system on the recently proposed MMHal-Bench [30], which involves 96 questions of a variety of types to which short answers are often appropriate, such as “How would you describe the weather in the image?". The corresponding images are hand-selected from OpenImages. The evaluation process sends ground truth annotations and the model’s responses to GPT-4, and requests GPT-4 to judge the descriptiveness and level of hallucinations. This is combined into a composite score (higher is better).
We show the results in Table 2, consisting of the average score as reported by the MMHal-Bench. We see that our approach results in a much larger gain in performance (+0.3) compared to the baseline established by LLaVA-7B, which is more substantial than the improvement achieved through the RLHF scheme proposed by LLaVA-RLHF [30] without using supervised fine-tuning (+0.1). We further note that our particular implementation of the ViGoR framework only uses prompts to elicit detailed descriptions of the entire scene, and does not cover additional types of questions. Therefore, the encouraging results on MMHal-Bench suggests that the improvements generalize well to different types of queries. As the underlying images do not come from MS COCO or ADE20K (the data sources for our fine-tuning process), this further shows the generalization capability of our approach.
4.4.3 Short Answer Programmatic Benchmark Evaluation.
Finally, we evaluate the models using benchmarks which consist of questions with concise and unambiguous one word answers such as “Is there a red apple in the image?" These benchmarks avoid the uncertainty and noise introduced from leveraging a third-party large language model for evaluation, and instead uses programmatic comparisons. Similar to MMHal-Bench, these benchmarks can measure both the visual grounding capabilities and the generalization of our method as the question types are outside the scope of the training data. We select the commonly used POPE [20] and MME [12] benchmarks. POPE probes the LVLM with 3000 yes/no questions targeting 500 images from COCO [21] to evaluate the LVLM’s ability to determine the existence of specific objects in a scene. MME extends the questions to cover aspects such as numerical count, position, and color of objects, as well as other vision tasks such as scene/landmark identification and OCR from data sources other than MS COCO and ADE20K.
Method #P VQA Random Popular Adversarial Avg F1
@@ 4.4 Quantitative Results
mPLUG-Owl 7B 53.3 51.7 50.6 50.3 50.7 50.3 67.2
MiniGPT-4 7B 77.8 75.4 68.3 64.3 66.6 62.45 74.1
MM-GPT 9B 50.0 50.0 50.0 50.0 50.0 50.0 70.1
InstructBLIP 13B 88.7 85.0 81.4 75.1 74.4 67.7 83.7
LLaVA 7B 54.4 52.3 52.4 51.3 50.8 50.4 67.8
LLaVA-RLHF 7B - - - - - - 78.2
LLaVA-RLHF 7B - - - - - - 82.7
ViGoR (ours) 7B 85.1 89.0 81.5 83.0 75.5 73.8 83.8
Table 3: Quantitative results on POPE benchmark. POPE is comprised of three parts, each generated by different sampling strategies: Random, Popular, and Adversarial Sampling. We report the Accuracy(AC) and Precision(PR) for each part, alongside the average F1 score (Avg F1) across all three parts.
Method #Param Data VQA EX CT PO CO PT CE SC LM AT OC Overall
@@
[…]
@@
Table 4: Quantitative results on MME benchmark. ‘VQA’ indicates whether the model is fine-tuned using the VQA datasets. Abbreviations explained: ‘EX’: Existence; ‘CT’: Count; ‘PO’: Position; ‘CO’: Color; ‘PT’: Poster; ‘CE’: Celebrity; ‘SC’: Scene; ‘LM’: Landmark; ‘AT’: Artwork; ’OC’: OCR. For a fair comparison, we compare the Large Vision Language Model using the Vicuna backbone.
$\dagger$\$$\dagger$\$footnotetext: We report the number of LLaVA v1.0 reproduced by ourselves.
As illustrated in Tables 3 and  4, our model shows consistent improvement across all categories evaluated in both datasets. Furthermore, while the competitive baseline methods [38, 10, 35, 17] use VQA datasets containing ground truth for these types of straightforward questions, our method is able to achieve comparable or superior results without fine-tuning on such resources. This suggests that the visual grounding abilities learned by the model through our process is general. When the model learns to generate accurate comprehensive descriptions for images, it is able to apply the capabilities to related reasoning tasks, suggesting that the logical deduction capabilities of the underlying pre-trained LLM is retained and leveraged. Furthermore, our 7B parameter model shows on par performance with some models with nearly double the number of parameters (13B). This underscores the efficiency and robustness of our model in handling diverse visual-language tasks. As with the case of MMHal-Bench, the MME benchmark draws imagery from sources other than MS COCO to encompass posters, celebrity images, scenes, landmarks, artwork, and OCR. Our method (using ADE20K and MS COCO in the pipeline) still shows significant improvements in these categories, further proving its generalizability.
In Table 3, we also compare our method with LLaVA-RLHF [30]. The results indicate that our method outperforms LLaVA-RLHF, even when the latter is trained with additional VQA data. This observation further demonstrates the superior visual grounding capabilities of our model, attributed to our fine-grained reward modeling strategy.
@@ 4.5 Qualitative Results
In Figure 3, we show qualitative results of our technique in action, focusing on a comparative analysis between our model and its starting point (LLaVA). This comparison is particularly relevant as our model undergoes fine-tuning based on the LLaVA model. By juxtaposing these two models, we highlight the enhancements and advancements our model has achieved, offering valuable insights into the efficacy and impact of our training strategies. More results are provided in the Supplementary Materials.
4.6 Fine-grained vs Holistic Evaluation by Reward Model
Reward model design MME
none 960.1
holistic-based 1027.4
fine-grained-based 1309.3
@@ 4.6 Fine-grained vs Holistic Evaluation by Reward Model
In this section, we compare the fine-grained analysis of sentences by the reward model with the more commonly used holistic evaluation of complete responses. This assessment is carried out in the context of reward modeling with human feedback. For the holistic reward model, we retrained a reward model using a singular, unified reward for each description. Specifically, holistic reward is determined based on the content of the entire description, which is negative if the description contains at least one erroneous sentence, and positive otherwise. As shown in Table 5, the fine-grain-based approach demonstrates better performance than the holistic approach, which shows that denser and more informative signals allow the reward model to better discern the quality of the generated text by creating more direct links between the visual features in the image and the sentences describing them.
5 Conclusion, Limitations, and Future Work
@@
We note that despite its successes, ViGoR nonetheless exhibits several limitations. The automated component of our reward generation relies heavily on the capabilities of the perception model (i.e. GroundingDINO). As such, current version is limited to objects suitable for detectors while not being applicable to stuff regions, attributes, or layouts. Furthermore, the human evaluation data used for training the reward model is specific to a particular LVLM’s architecture and checkpoint with which the evaluated responses are generated. For LVLMs with significantly different output or failure modes, additional human preference data may need to be collected, thus incurring further cost.
We hope to address the aforementioned limitations in future work. As well, we plan to apply RLHF training with our ViGoR framework, which may offer further improvements over our training based on rejection sampling. As well, we anticipate that explicitly linking visual entities with associated phrases in the generated text can further improve visual grounding.